### v1

# Assist Research

Abstract
Numerous studies have assessed the proficiency of AI systems, particularly large language models (LLMs), in facilitating everyday tasks such as email writing, question answering, and creative content generation. However, researchers face unique challenges and opportunities in leveraging LLMs for their own work, such as brainstorming research ideas, designing experiments, and writing or reviewing papers. In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in four fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not. AAAR-1.0 differs from prior benchmarks in two key ways: first, it is explicitly research-oriented, with tasks requiring deep domain expertise; second, it is researcher-oriented, mirroring the primary activities that researchers engage in on a daily basis. An evaluation of both open-source and closed-source LLMs reveals their potential as well as limitations in conducting sophisticated research tasks. We will release the AAAR-1.0 and keep iterating it to new versions. Project Webpage:   https://renzelou.github.io/AAAR-1.0/
Figure 1: The input-output illustration of four tasks in the proposed AAAR-1.0 benchmark.
1 Introduction
Although AI has brought transformative changes to various aspects of life, its impact on researchers unfolds in a nuanced manner. On the one hand, AI assists in various research disciplines, such as Social Science (Neuman et al., 2023), Finance (Gu et al., 2024), Medicine (Rakhimov et al., 2022), GeoScience (Praskievicz, 2018), etc., significantly expediting academic processes. However, many of these applications are superficial, often limited to data-driven clustering or classification. On the flip side, the AI era poses challenges for researchers. Despite its ability to streamline some activities, researchers still face demanding, cognitively intensive tasks such as staying current through extensive paper reading, rapidly generating ideas in response to fast-paced advancements, conducting rigorous experiments to substantiate claims, and managing an increasing volume of peer reviews. Then a question looms: How effectively can AI assist researchers in tasks that are domain-specific, expertise-demanding, and reasoning-intensive?
Existing works proved the promising potential for using LLMs in assisting AI research. Si et al. (2024) conducted a large-scale human study and found that LLMs can generate creative research ideas. Lu et al. (2024) proposed an autonomous agent to handle complicated research workflow and write a whole research paper. However, most of these works focus on addressing highly subjective problems that require a high degree of expertise, making evaluation laborious and hard to reproduce. This underscores the need for a comprehensive benchmark that rigorously assesses LLMs’ capabilities in expertise-intensive research activities.
To this end, in this work, we introduce AAAR-1.0, a novel benchmark that aims to comprehensively assess the LLMs’ capacity on expert-level research tasks. As illustrated in Figure 1, AAAR-1.0 decomposes four distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews. To ensure data quality, senior AI researchers with extensive domain expertise perform data annotation for AAAR-1.0, followed by rigorous multi-round data examination and filtering. All four tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed. Crucially, tasks here are singular, stand-alone challenges (with clear input and output expectations) rather than a complicated task chain (Li et al., 2024; Lu et al., 2024), providing a more transparent assessment of the model’s intermediate output.
Benefiting from the proposed automatic metrics, we conduct extensive experiments across numerous mainstream LLMs, where we find that:
  • With a random guess baseline of 25%, the performance of most LLMs on EqInfer hovers just slightly above chance, with the top models reaching around 60%. This highlights the difficulty of the task, despite its reliance only on local context reasoning.
  • In ExpDesign, LLM-designed experiments are innovative and more diverse than those by humans; however, many are trivial, lack feasibility, and stray from the original research objectives.
  • In PaperWeakness, LLM-identified weaknesses often lack depth and specificity, making them broadly applicable and less useful for providing feedback on paper drafts.
  • In ReviewCritique, LLMs struggle to effectively identify deficient human reviews, indicating limited usefulness in assisting meta-reviewers in evaluating the quality of individual paper reviews.
2 Related Work
LLMs for AI Research.
With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024). For example, Si et al. (2024) conducted a large-scale human investigation on LLM-generated research ideas and found that LLMs can generate novel ideas compared with humans while lacking feasibility. Du et al. (2024) found that while LLMs are effective at summarizing papers, they tend to overly trust the authors’ claimed strengths and struggle to identify weaknesses specific to the paper. Furthermore, some works try to employ LLMs to solve more complicated research tasks that are composed of multiple steps (Li et al., 2024; 2023; Tang et al., 2023). Notably, Lu et al. (2024) proposed AI-Scientist, an autonomous agent framework that can handle a series of challenging research tasks consecutively, including generating research ideas, coming up with the corresponding experiments along with the implementations, and then writing the final research paper — exactly how human conduct a whole research pipeline. However, there is still a lack of systematic evaluations and quantitative analyses on the LLMs’ (intermediate) output of each single-step research task. Our work focuses on building a benchmark that has individual research steps with clear input-output expectations, thus making it suitable for comprehensive LLMs evaluation.
Benchmarks for AI Research Tasks.
Existing “LLM assists research” benchmarks mainly focus on the implementation and execution part of the research pipeline (Lu et al., 2024; Chen et al., 2024b; Li et al., 2024). For instance, Huang et al. (2024) proposed MLAgentBench to test the LLMs’ capacity for writing project code and training the ML models, where the evaluation metric is the test performance of the models trained by LLMs. However, real-world AI research activities are diverse and some of them are hard to assess for quality, such as generating research ideas, which requires intensive manual assessment (Si et al., 2024; Liang et al., 2024), or LLM-based estimation (Lu et al., 2024). Our work centers on tasks that emphasize a comprehensive mastery of the scientific research field and core elements of a researcher’s daily workload, and we try to build curated task-specific metrics for every single task for a more efficient and accurate LLMs appraisal.
Figure 2: Data construction workflows of the three tasks in Refer to caption AAAR-1.0.
3 AAAR-1.0
Figure 2 provides an overview of the construction process of AAAR-1.0. In the following sections, we elaborate on the data collection details of the aforementioned four tasks, including § 3.1 EquationInference (EqInfer), § 3.2 ExperimentDesign (ExpDesign), § 3.3 PaperWeakness (Weakness), and § 3.4 ReviewCritique.
3.1 EquationInference
Crafting a correct scientific equation in paper writing or validating an equation in paper reviewing is challenging, as it requires a thorough understanding of an algorithm or the intricate relationships among numerous variables. Directly prompting LLMs to generate equations proves overly demanding. Therefore, this work formulates EqInfer (Figure 1) as as a multiple-choice equation inference task based on contextual cues, curated as follows.
① Data crawling and cleaning.
For the data source, we adopt the pre-compilation LaTeX code for two reasons: i) existing PDF parsing tools, such as PyMuPDF and PaperMage (Lo et al., 2023), can introduce considerable noise to the parsed equation text; ii) considering most of exiting LLMs are capable with processing LaTeX code, using LaTeX source instead of parsed text can be more accurate and provide LLMs with richer information. Meanwhile, to avoid using any low-quality human-written equations, we only crawl those peer-reviewed papers accepted by top-tier conferences. Accordingly, we first obtain the accepted paper list from ACL Anthology, from year 2019 to 2023. Next, we search each paper on arXiv to crawl its LaTeX source (if it exists). Finally, we get a total of 1,762 papers’ source LaTeX packages.
We then clean the LaTeX sources by deleting all the comments and combining multiple cross-referred .tex files into a main file. Afterwards, we use regex to randomly extract (at most) 3 equations’ code snippets per paper, finally resulting in 3,877 human-written equations are extracted.
② LLM-based equation synthesis.
As we formulate this task as classification, for each human-written positive equation, we have to craft at least three counterpart negative equations. To this end, we prompt GPT-4 to synthesize more equations based on the paper context. For each positive equation, we repeat this prompting (with a high decoding temperature) until three different negative equations are synthesized.
③ LLM-based filtering.
However, the LLM-synthetic equations can sometimes be context-unaligned, i.e., some synthesized equations contain notations that are never defined in the paper context, which is a superficial shortcut for the classification tasks. To improve the data quality, we prompt GPT-4 to identify those context-unaligned negative equations. We then discard those instances where all three negative equations are identified as contextually unaligned. This filtering leads to a final of 1,449 classification instances (62.3% instances are filtered).
④ Expert-based examination.
Furthermore, it’s also possible that synthesized negative equations are actually correct (i.e., false negative options) — even if the negative and positive equations are written differently, the final compiled results might be the same. To filter out the false negative equations and to have a final check on the classification instances, we then employ human experts to conduct a further data review.
We asked 5 senior Ph.D. students who are experienced in AI research to manually check all the instances. For each classification instance, we ask those human experts to consider the following criteria: i) Are all four equations (both positive and negative) grammatically correct? ii) After compilation, is there only one correct answer? We ask every human expert to use external LaTeX compilation tools (e.g., TeXlive), and identify the instances that cannot meet the criteria. Each instance is examined by at least two experts, and we only keep instances that all experts decide to keep. After this strict examination, a total of 1,049 instances are eventually kept (27.6% instances are filtered)
Final data.
We finally shuffle the four equations for each classification instance and randomly assign letters (A, B, C, and D) to the equations. We show the data statistics of the final EqInfer in Table 11 and the sample data cases in Figure 7.
3.2 ExperimentDesign
Given a research topic, such as a novel ML algorithm, a qualified researcher can design a solid experiment plan for it, and clarify underlying motivation to ensure the reliability of the designed experiment. Unlike the concurrent works that focus on the experiment implementation (Lu et al., 2024; Huang et al., 2024), we emphasize the importance of assessing the high-level experiment design of LLMs before the subsequent implementation to avoid any expensive execution iteration. Therefore, as shown in Figure 1, we formulate ExpDesign as a text-generation task that takes pre-experiment paper context as input, and then generates the experiment and explanation list.
① Data crawling.
As for the data source, we first collect $\geq$ 10k papers’ data from arXiv, including LaTeX sources and PDFs, which cover broad AI categories, including cs.AI, cs.CL, and cs.CV, from year 2018 to 2023. Similarly, to ensure the source data quality, we only use papers that have appeared at well-known conferences.
② Domain-expert annotation.
Making a reliable and executable experiment plan requires solid foundation knowledge of a specific research area. Consequently, we set a high standard for choosing annotators: i) be a senior Ph.D. student with at least one peer-reviewed publication in leading AI venues; ii) have more than 4 years of AI research experience; iii) frequently serve as conference reviewers. Finally, we invite a total of 10 qualified experts to participate in our data collection procedure. Given the 10k crawled papers, we first ask every annotator to bid on the papers that they are interested in. After bidding, each of them is assigned 10 papers, i.e., a total of 100 papers to be annotated. During annotation, we post each paper PDF on online Google Drive and ask the annotator to first carefully read the whole paper. Then, we ask them to identify and locate the key experiments in each paper (i.e., highlighting the relevant paragraphs of each experiment). We don’t consider some trivial experiments, such as those supplemental analyses in the appendix section. For each identified experiment, the annotator has to concisely answer two questions: i) What did this experiment do? ii) Why did the paper authors conduct this experiment? In other words, we ask the annotator to summarize all the key experiments in this paper and explain the underlying motivations based on their rich domain experience.
③ Multi-round peer discussion.
Intuitively, different experts might have different opinions on the same research topic. Particularly, when explaining the underlying motivation of an experiment, adopting only a single expert’s opinion might introduce bias to our annotation. Hence, we conduct a further multi-round peer discussion. For each online paper PDF, where all the key experiments are identified, summarized, and explained, we ask a different expert (reviewer) to review the annotation by considering the following three criteria: i) Are the identified experiments all the key experiments? ii) Does each experiment summarization covers all key information? iii) Does each explanation sound reasonable and reliable? Each reviewer has to leave comments to the online PDF regarding the above criteria, and then the annotator has to respond to each comment — either accept the suggestion and revise the previous annotation, or provide a “rebuttal” to the reviewer to uphold the annotation. This discussion iterates until both opinions align with each other. Eventually, for each paper, we collect two lists: i) the experiment list, summarizing each experiment step of the paper; ii) the explanation list, the underlying motivations that are one-one corresponding to the experiment.
Final data.
After annotation, we use the pre-experiment context of each paper (according to the first-experiment location identified by the annotator) as the input. Furthermore, we use GPT-4 to delete any sentence that potentially leaks the experiment from the input. [^1] About 9.8% sentences are deleted. Similar to the EqInfer, we utilize the source LaTeX as the input text to avoid PDF paring noise. As for the image input, we collect those figures within each paper’s source LaTeX package and only keep figures that are used in the pre-experiment context. Overall, a total of 100 instances are collected. As shown in Figure 1, the input of each instance is the pre-experiment context (including the figures), and the ground-truth output is the expert-annotated experiment plan and the explanations. Table 12 shows data statistics and Figure 8 illustrates the sample case in ExpDesign.
3.3 PaperWeakness
Another critical research task is paper review. Previous works have demonstrated the usefulness of the LLM-based review feedback (Gao et al., 2024; Jin et al., 2024; Lu et al., 2024). However, as indicated by Du et al. (2024); Liang et al. (2024), LLMs only excel at summarizing the research strengths while falling significantly short on weakness criticism. Hence, we build Weakness for particularly investigating the LLM-generated weaknesses.
① Data crawling.
We first crawl a total of 3,779 anonymous submissions of ICLR 2023 from OpenReview, [^2] We adopt ICLR because it releases full submissions, while some other conferences only release accepted papers. including PDF and other meta information (e.g., scores, decisions, and tracks). As the ICLR 2023 has 13 distinct tracks while the paper distribution across different tracks is highly biased, we then uniformly sample papers from different research tracks to improve the domain diversity. Meanwhile, during sampling, we also keep the accept/reject papers distributed equally to avoid data bias. In a word, we finally collect a total of 1,000 papers (500 accepted; 500 rejected), uniformly covering all 13 tracks. Please refer to Figure 5 for the track and score distribution of the 1,000 papers.
② Extraction of human-written weaknesses.
Since the raw comments crawled from ICLR 2023 are mixed with both strengths and weaknesses, we further employ GPT-4 to extract all the weaknesses from each reviewer’s comments and compose multiple weaknesses into a list. Notably, we force GPT-4 to keep the original text of the reviewer, i.e., all weaknesses in our dataset are those original sentences written by the reviewer without any modifications. [^3] We manually checked GPT-4’s extraction results of 200 cases — GPT-4 only missed $\leq$1of reviewer-written weaknesses and maintained almost all the original text. What’s more, sometimes one reviewer might repeatedly mention the same weakness throughout the comment. In this case, we simply keep all the repeated weaknesses because, if one weakness is repeatedly mentioned by the reviewer, it’s intuitively an important weakness that the reviewer wants to emphasise; accordingly, keeping the repeat items can penalize LLMs more on missing this weakness.
For each paper, we can finally get multiple weakness lists (one weakness list per reviewer, one paper can have multiple reviewers). We further delete a few papers without any weaknesses found in the raw comments, resulting in a total of 993 instances, i.e., 993 {paper, weakness lists} pairs.
③ Input data processing.
As we mentioned before, we crawl papers from OpenReview instead of arXiv because the under-review paper draft is required for this task. However, not every paper from OpenReview can be found on arXiv, i.e., the source LaTeX code and figures of most under-review papers are unavailable. Therefore, we utilize VILA (Lin et al., 2023) to parse text data out from the PDF; we also employ PDFFigures-2.0 (Clark & Divvala, 2016) to extract all the figures and tables (in image) from the paper, as Vila is not good at processing the table data.
Final data.
Our final data is composed of 993 instances, each input is paper text along with figure/table images, and each output is peer reviewers’ weakness lists. Table 13 shows data statistics; Figure 9 presents an example of the data instances. We show the data diversity (score and track distribution) in Figure 5.
3.4 ReviewCritique
In addition to identifying weaknesses in paper drafts, a more challenging research task that requires more senior research experience is conducting meta-reviewing. Given a paper submission, along with individual reviews and author rebuttals, meta-reviewing is not to summarize individual reviews. Instead, a meta-reviewer must go through all the information and make a final recommendation. This requires the meta-reviewer to identify deficient/unreliable review segments (e.g., if a viewpoint is too subjective or contains factual errors) in each individual review and make a decision based on the non-deficient ones. This task demands years of experience in the relevant domain; even for human experts, only senior researchers are typically qualified for meta-reviewing. Therefore, as illustrated in Figure 1, we also investigate how LLMs assist meta-reviewers, specifically in identifying deficient review points.
We reuse the ReviewCritique dataset from our recent work (Du et al., 2024), where we crawled papers’ initial submissions along with their reviews from OpenReview and employed more than 40 AI research experts to label each review segment (i.e., deficient or not), with detailed human explanations. In total, there were 100 papers with 380 human reviews. Each review was divided into sentence-level segments, resulting in 11,376 review segments (viewpoints).
4 Evaluation Criteria
For EqInfer, we adopt accuracy as the classification criterion. For ExpDesign and Weakness, since both tasks have natural language outputs, semantic-based metrics are necessary. Hence, in addition to the conventional ROUGE (Lin, 2004), we also develop several novel similarity-based metrics for each specific task.
When evaluating the experiment plan list of ExpDesign, we hope the LLMs can mention as many similar experiment steps as the expert’s plan. Nevertheless, we also don’t expect LLMs to generate too many irrelevant or redundant steps in the plan. This intuition covers both the “recall” and “precision” aspects. Therefore, we develop semantic similarity-based F1 score, denoted as S-F1, which is the harmonic mean of S-Precision and S-Recall:
S-Precision $=\frac{1}{m}\sum_{i=1}^{m}\max_{j}\textrm{sim}(p_{i},g_{j})$ (1)
S-Recall $=\frac{1}{n}\sum_{j=1}^{n}\max_{i}\textrm{sim}(g_{j},p_{i})$ (2)
where the $p$ and $g$ represent the LLM’s prediction plan and the ground-truth plan, respectively. The $m$ and $n$ are the list length of $p$ and $g$ (e.g., $m$ experiment steps in $p$). We use SentenceBERT (Reimers, 2019) to measure the semantic similarity between the $p_{i}$ step and the $g_{j}$ step.
Meanwhile, S-F1 omits the item order difference of two lists, but when giving same-length lists (items have one-one correspondence), we can utilize the following similarity-based matching score:
[…]
Unlike ExpDesign, the output of Weakness is multiple reviewers’ weakness lists, which means we have to measure LLM’s single prediction list with a “nested” list. Hence, we rewrite S-Precision ,  S-Recall to SN-Precision , SN-Recall:
SN-Precision $=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{r}\sum_{k=1}^{r}\max_{j}\textrm{sim}(p_{i},g^{k}_{j})\right)$ (4)
SN-Recall $=\frac{1}{r}\sum_{k=1}^{r}\left(\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\max_{i}\textrm{sim}(g^{k}_{j},p_{i})\right)$ (5)
where $r$ is the number of reviewers of the given paper, $n_{k}$ means the length of $k$-th reviewer’s weakness list, and $g^{k}_{j}$ indicates the $j$-th item in $k$-th reviewer’s weakness list.
Additionally, in the real world, we would think a review weakness is reliable if it is specific to a paper. Meanwhile, we also hope the review is informative, i.e., no excessive similar weaknesses in one review. Therefore, we adopt the following ITF-IDF metric proposed by Du et al. (2024), which is inspired by the classic TF-IDF:
ITF-IDF $=\frac{1}{w}\sum_{j=1}^{w}\left(\frac{1}{m_{j}}\sum_{i=1}^{m_{j}}\log\left(\frac{m_{j}}{O_{i}^{j}}\right)\times\log\left(\frac{w}{R_{i}^{j}}\right)\right)$ (6)
$O_{i}^{j}$ $=\sum_{k=1}^{m_{j}}\textrm{sim}(p_{i}^{j},p_{k}^{j})$ (7)
$R_{i}^{j}$ $=\sum_{l=1}^{w}\max_{s}\textrm{sim}(p_{i}^{j},p_{s}^{l})$ (8)
where the $w$ is the total number of papers in the dataset, $p^{j}$ is $j$-th paper’s prediction weakness list, $p^{j}_{i}$ is the $i$-th weakness in $p^{j}$. Moreover, ${O_{i}^{j}}$ calculates the intra-paper occurrence frequency of $p^{j}_{i}$; ${R_{i}^{j}}$ is the “soft” number of papers that also contain the $p^{j}_{i}$, which is computed by summing the maximum similarity scores between $p^{j}_{i}$ and other paper’s weaknesses. In a word, ${O_{i}^{j}}$ measures informativeness, and ${R_{i}^{j}}$ measures specificity. The complete ITF-IDF consider both aspects and reflects the overall weakness diversity.
For ReviewCritique, we use F1 score as the classification metric; while for the deficiency explanation, we use ROUGE (Lin, 2004) and BERTScore (Zhang et al., 2020) to reflect how well the model-generated explanation aligns with the expert’s annotation.
5 Experiments and Analyses
In this section, we conduct extensive experiments on AAAR-1.0, across various mainstream LLMs, to quantify the current LLMs’ capacity to tackle high-level research tasks. Specifically, § 5.1 for EqInfer, § 5.2 for ExpDesign, § 5.3 for Weakness, and § 5.4 for ReviewCritique. Please refer to Appendix B.2 for running details of the LLMs.
5.1 EquationInference
Table 1: Various LLMs’ performances on the 1,049 instances of EqInfer task.
Methods Accuracy (%)
Random Guess 25.00
Open-source LLMs
Gemma 2-27B (Gemma Team, 2024) 3.24
Falcon-40B (Almazrouei et al., 2023) 4.39
OLMo-7B (Groeneveld et al., 2024) 19.00
Mistral-7B (Jiang et al., 2023) 22.21
Qwen 2.5-72B (Qwen Team, 2024) 35.93
Mixtral-8x22B-MoE (Jiang et al., 2024) 37.08
Llama 3.1-70B (MetaAI, 2024) 38.13
Closed-source LLMs
Gemini 1.5 Pro (Anil et al., 2023) 34.31
GPT-4o (OpenAI, 2024a) 43.18
GPT-4 (OpenAI et al., 2023) 49.85
o1-preview (OpenAI, 2024b) 59.49
Claude 3.5 sonnet (Anthropic, 2024a) 61.10
Figure 3: The input context length scaling trend on the EqInfer task.
Settings.
As different LLMs have distinct context windows, to ensure a fair comparison, we fix the maximum input length for all models. According to the data statistics of Table 11, we empirically use 1,000 words for both contexts before and after equations, i.e., 2,000 surrounded words.
Main results.
Table 1 shows the main results. Firstly, the open-source LLMs, especially the Falcon and Gemma, perform unexpectedly disappointing (even worse than random guesses). These screwed scores are mainly due to the poor long-context instruction following ability, where we find some open-source LLMs are confused with the massive input and often copy the LaTeX code from the input. In contrast, closed-source LLMs generally achieve superior accuracy, probably owing to the richer scientific knowledge from the larger model parameters. However, considering the conventional multi-choice QA formulation of EqInfer, the recently-released GPT-4o solely gets 43.18, implying the unique challenge of EqInfer compared with other scientific QA benchmarks (Song et al., 2023). Notably, with the help of internal CoT, o1 gains stronger performances than GPT-4/GPT-4o, indicating the potential benefits of adopting reasoning for this task.
$\mathcal{Q}$: Do more contexts boost performance?
Table 1 unifies the input context lengths to 1,000 tokens for various LLMs. To answer this question, we experiment with long-context LLMs to investigate the impact of the input context lengths. Particularly, we scale the input length (per side) from 100 to 1,500 words. As shown in Figure 3, for the open-source LLMs (Llama and Qwen), after 300 words length, increasing the input context doesn’t help the performance and even significantly drops Qwen’s scores. While for the closed-source GPT-4-Turbo and GPT-4o, scaling up input length gradually boosts the performances at the first 1,000 words, but stabilizes afterwards. This is in line with human intuition, i.e., surrounding context is required for the equation inference, as the adjacent context usually provides important information, such as the target algorithm description or the notation definition. However, after exceeding a specific threshold, more context information is not beneficial anymore and even confuses those LLMs with poor long-context handling capacity (Wang et al., 2024; Liu et al., 2024).
5.2 ExperimentDesign
Table 2: Various LLMs’ performances on the 100 instances of ExpDesign. The explanation generation is based on the oracle experiments to prevent error propagation. “Copy Input” is a random baseline: for experiment design, randomly select 5 sentences from the input paper; for experiment explanation, directly copy each experiment idea.
Experiment Design Experiment Explanation
Methods S-F1 S-Precision S-Recall S-Match ROUGE-L ROUGE-1
Copy Input 21.13 17.94 26.76 40.32 22.06 25.28
Open-source LLMs
OLMo-7B (Groeneveld et al., 2024) 33.94 37.25 31.79 45.78 26.30 30.38
Falcon-40B (Almazrouei et al., 2023) 17.87 21.78 15.35 17.03 12.10 12.72
Gemma 2-27B (Gemma Team, 2024) 34.33 39.71 30.51 42.77 26.20 29.63
Mistral-7B (Jiang et al., 2023) 37.62 43.09 34.19 50.18 30.20 34.69
Mixtral-8x22B-MoE (Jiang et al., 2024) 42.21 50.13 36.82 49.07 29.96 34.53
Llama 3.1-70B (MetaAI, 2024) 40.57 48.43 35.43 50.05 29.33 34.11
Qwen 2.5-72B (Qwen Team, 2024) 43.24 51.73 37.55 51.12 29.46 34.68
Closed-source LLMs
Gemini 1.5 Pro (Anil et al., 2023) 51.87 50.77 53.37 52.87 28.52 33.80
Claude 3.5 sonnet (Anthropic, 2024a) 48.74 46.49 51.53 53.03 18.75 26.15
GPT-4 (OpenAI et al., 2023) 43.89 42.34 45.82 55.03 22.82 30.01
GPT-4o (OpenAI, 2024a) 53.00 51.24 55.12 54.79 27.54 34.31
o1-preview (OpenAI, 2024b) 46.67 45.04 48.70 58.55 29.11 36.70
Settings.
Similarly, we unify the input context length of different LLMs to ensure a fair comparison. According to Table 12, we set 2,000 and 3,000 input words for open- and closed-source LLMs, respectively. Meanwhile, as experiment explanation is the subsequent task of experiment design, using model-generated experiments can propagate errors in explanation, leading to inferior results for most LLMs. To this end, we provide LLMs with the oracle experiments when generating explanations.
Main results.
Table 2 shows the main results. For the experiment design, the closed-source LLMs generally outperform open-source LLMs, and both closed-/open-source LLMs are superior to the “Copy Input” baseline (except the Falcon). Despite the higher S-Precision, the open-source LLMs are seriously deficient in S-Recall compared with closed-source LLMs ($\sim$10$\downarrow$). We find that closed-source LLMs are more creative in experiment design and tend to generate more experiment ideas than open-source LLMs (though most of the experiment ideas are trivial), leading to excellent S-Recall. As for the experiment explanation, the S-Match scores of closed-source LLMs still surpass the open-source LLMs, while the score difference is not significant. Furthermore, we find the negative correlation between S-Match and the ROUGE, where the ROUGE scores of closed-source LLMs are broadly inferior. We find that the open-source LLMs often try to copy the terms or phrases from the given experiment, or even simply paraphrase the experiment instead of explaining, which results in a high superficial overlap with the ground-truth explanation. This observation highlights the importance of adopting the proposed S-Match to avoid evaluation bias of traditional generation metrics.
$\mathcal{Q}_{1}$: Can self-contained experiment design enhance the experiment explanation?
Table 3: The impact on S-Match scores of maintaining the experiment’s self-containment for ExpDesign.
Models One-by-One Whole-List
Llama 3.1-70B 50.05 49.36 ($\downarrow$ 0.7)
Qwen 2.5-72B 51.12 48.56 ($\downarrow$ 2.6)
Gemini 1.5 Pro 52.87 57.48 ($\uparrow$ 4.6)
Claude 3.5 sonnet 53.03 59.11 ($\uparrow$ 6.1)
GPT-4 55.03 56.95 ($\uparrow$ 1.9)
GPT-4o 54.79 58.54 ($\uparrow$ 3.8)
o1-preview 58.55 61.58 ($\uparrow$ 3.0)
When generating the explanation in Table 2, we provide LLMs with each individual experiment and let them explain one by one, because we find that, when providing the whole experiment list, those open-source models only explain partial experiments because of their poor instruction-following capacity. However, there are intuitively some semantic or logical relations between different experiments, e.g., some experiments are prerequisites to others. Therefore, this one-by-one prompting might break the self-containment of an experiment plan. Consequently, we test with the “whole-list” prompting, where the LLMs are given the complete experiment list and are asked to explain all experiment steps together.
As shown in Table 3, unlike the open-source LLMs, the explanation performances of those closed-source LLMs are generally improved after adopting whole-list prompting. According to further manual checking, after maintaining the self-containment of the experiments, the LLMs can refer to other experiments and better grasp the underlying motivation of the current experiment.
$\mathcal{Q}_{2}$: Do human evaluation results align with automatic metrics for explanation?
Table 4: The human evaluation results on LLMs’ output explanations of ExpDesign. “Acc. ratio” means how many model outputs are accepted by the annotator.
Models Acc. ratio
Llama 3.1-70B 22.93
Gemini 1.5 Pro 55.07
Claude 3.5 sonnet 61.46
GPT-4o 69.72
o1-preview 76.14
As the explanation can be open-ended, in this paragraph, we provide the human evaluation results on different LLMs’ experiment explanation outputs. In detail, we randomly select 20 out of 100 papers and ask 5 annotators to read the experiments along with each model’s explanations; we then let the annotator decide whether each model’s explanation is acceptable (see Appendix C.1 for more details). Table 4 illustrates the results, where the score variance is higher than Table 2. However, the performance ranking of both tables is perfectly correlated with each other (Spearman’s rank correlation coefficient $=$ 1), demonstrating the effectiveness of S-Match.
$\mathcal{Q}_{3}$: Do more contexts boost performance?
Figure 4: The input context length scaling trend of different LLMs on the ExpDesign task.
We also investigate the impact of input context length for ExpDesign. As shown in Figure 4, we scale up the input pre-experiment context length from 0.1k to 10k tokens (10k tokens is the maximum paper context length in the dataset). For the experiment planning, more input context does improve the performance of different LLMs, while this benefit stops after exceeding 5k tokens, which is similar to EqInfer’s scaling results — after the necessary information has been covered, scaling more up doesn’t boost the performance. Meanwhile, the results of the experiment explanation demonstrate that explaining motivations almost doesn’t require any paper context, i.e., the LLMs solely rely on the given experiments. However, we do not expect this because we hope LLMs can explain the motivation based on a thorough understanding of the paper, just like how human experts do. Hence, there is still a considerable gap between the LLMs and humans in terms of grasping research motivations.
$\mathcal{Q}_{4}$: Does multi-modal input boost performance?
Table 5: The figure inputs ablation of ExpDesign. For the maximum text input length, same as the setting in Table 2, we use 2,000 and 3,000 words for open- and closed-source models, respectively. For the closed-source GPT-4o and GPT-4, as they have long context window sizes, we use all the figures of each paper. While for InternVL2, we randomly select two figures per input paper.
Experiment Design Experiment Explanation
Models S-F1 S-Precision S-Recall S-Match ROUGE-L ROUGE-1
GPT-4o 53.00 51.24 55.12 58.54 29.25 35.50
w/ figures 50.11 48.94 51.59 58.53 27.87 34.30
GPT-4 43.89 42.34 45.82 56.95 25.98 33.37
w/ figures 43.54 42.56 44.85 55.03 22.82 30.01
InternVL2-26B 40.52 48.95 35.20 50.03 29.13 34.26
w/ figures 38.83 46.91 33.70 50.29 29.29 34.06
Intuitively, besides the text, when designing experiments for a given research topic, the figures can provide rich supplementary information, such as an algorithm illustration that can help better understand this research topic and underlying motivations. Hence, we test the performance of different LMMs (Large Multimodal Models), including GPT4-o, GPT-4, and InternVL2 (Chen et al., 2024a). Table 5 shows the ablation results on the figure data. To our surprise, the figure data doesn’t improve the LMMs’ results in this task, even harming the performances. This might be due to the low informativeness of the figures, as figures usually consume more input tokens but act only as supplementary information to the text, indicating future work on developing LMMs that can effectively leverage the scientific figures.
5.3 PaperWeakness
Settings.
Intuitively, the full paper content is necessary for providing feedback to that paper. Therefore, instead of setting a maximum input length, in Weakness, we try to feed all the paper context into the LLMs. As the input length of Weakness is extremely long (see Table 13), we adopt a “split-combine” method — we first split the whole paper into several smaller pieces and let LLMs predict the weaknesses of each piece separately; after that, we combine all pieces’ weaknesses as a final complete prediction. In practice, for the length of each small piece, we set 2,000 and 3,000 words for open- and closed-source LLMs, respectively. Additionally, in this task, we also examine the performance of a recent agent framework, namely AI-SCI (Lu et al., 2024), which enhances GPT-4o’s paper review ability by leveraging advanced prompting techniques, e.g., self-reflection (Shinn et al., 2024) and response ensembling (Wang et al., 2023). [^4] We don’t run AI-SCI on ExpDesign, because AI-SCI takes model-generated ideas as the inputs, which are incompatible with our task setting.
Main results.
Table 6: Various LLMs’ performances on the 993 instances of Weakness.
Weakness Diversity
Methods SN-F1 (%) SN-Precision (%) SN-Recall (%) ITF-IDF ($\uparrow$)
Human Review 7.69
Open-source LLMs
OLMo-7B (Groeneveld et al., 2024) 43.25 40.38 47.04 2.45
Falcon-40B (Almazrouei et al., 2023) 27.34 25.13 30.88 1.06
Gemma 2-27B (Gemma Team, 2024) 35.85 34.68 37.91 1.43
Mistral-7B (Jiang et al., 2023) 42.03 43.80 40.77 1.17
Mixtral-8x22B-MoE (Jiang et al., 2024) 43.23 44.59 42.23 0.98
Llama 3.1-70B (MetaAI, 2024) 42.78 43.19 42.70 2.60
Qwen 2.5-72B (Qwen Team, 2024) 42.74 43.80 42.05 1.21
Closed-source LLMs
Gemini 1.5 Pro (Anil et al., 2023) 48.75 43.97 55.08 5.88
Claude 3.5 sonnet (Anthropic, 2024a) 47.85 41.97 56.00 3.91
GPT-4 (OpenAI et al., 2023) 47.66 42.15 55.19 5.31
GPT-4o (OpenAI, 2024a) 47.73 42.09 55.48 5.95
o1-preview (OpenAI, 2024b) 48.62 42.54 57.08 5.63
LLM Agent Framework
AI-SCI (GPT-4o) (Lu et al., 2024) 45.05 40.02 51.91 2.23
Table 6 shows the main results, where the closed-source LLMs’ overall performances are generally superior to the results of open-source LLMs. Similarly, closed-source LLMs are particularly excellent in SN-Recall because of more generated weaknesses. However, there is still a considerable gap in the weakness diversity between the LLMs and human experts. [^5] Note that the human’s ITF-IDF score in Table 6 can be slightly underestimated. This is because we keep the repeated weaknesses in the human review, which affects the human review’s informativeness (lower ITF) but is useful when calculating the SN-Recall for LLMs. Compared with human review, most LLM-generated weaknesses are vague and lack the necessary knowledge about some frontier research works. Surprisingly, AI-SCI performs worse than backbone GPT-4o, especially on ITF-IDF, which suggests the challenge of Weakness, i.e., simply adopting popular prompting techniques cannot well address this task.
Table 7: The performance comparison of different input processing methods for Weakness. We use GPT-4o and GPT-4-Turbo because both accept a maximum of 128k tokens input. We also put the results of AI-SCI in the table for reference. Here, “split-combine” splits the input paper into several pieces, where each piece’s length is denoted as “window size”; “no-split” means the conventional input cutting, for example, if the window size is 3,000, then only the first 3,000 words in the paper are used. According to the data statistics, 20,000 words can cover maximum lengths of more than 95% of the papers in our dataset.
Models Input Context Processing Window Size (in words) SN-F1 SN-Precision SN-Recall ITF-IDF
GPT-4-Turbo split-combine 3,000 47.66 42.15 55.19 5.31
no-split 3,000 45.80 43.66 48.39 5.58
no-split 20,000 44.99 42.64 47.82 5.58
GPT-4o split-combine 3,000 47.73 42.09 55.48 5.95
no-split 3,000 45.74 43.45 48.54 5.92
no-split 20,000 45.47 42.97 48.51 6.02
AI-SCI split-combine 3,000 45.05 40.02 51.91 2.23
no-split 3,000 42.56 40.90 44.65 2.53
no-split 20,000 42.53 40.75 44.78 2.58
$\mathcal{Q}_{1}$: Is the split-combine effective?
Ideally, if the LLM has a sufficient context window size, it is unnecessary to split the input papers for separate processing. Consequently, in this paragraph, we utilize the LLMs accepting long context input to compare “split-combine” with “no-split”, i.e., letting LLMs write weaknesses by giving the full paper. In practice, we set the maximum number of input words to 20k, which ensures $\geq$95papers in the Weakness can be fully processed. As shown in Table 7, compared with giving the full paper contexts, split-combine generally brings about superior performances. During manual checking, we find that, when full paper is available, LLMs frequently neglect some important sections and omit weaknesses accordingly, while split-combine ensures that the LLMs can carefully brainstorm weaknesses within each smaller piece. Surprisingly, the LLMs’ performances with full paper context can be even worse than just remaining the first 3,000 words. This implies that even the current powerful long-context LLMs still fall short when processing long scientific documents.
$\mathcal{Q}_{2}$: Does multi-modal input boost performance?
Table 8: The ablation study about the paper tables and figures of Weakness. Based on the conclusion in Table 7, we use the “split-combine” to process the text input here (2,000 and 3,000 words context window size for open- and closed-source models). For GPT-4o, we use all the table/figure images; while for InternVL2, we randomly select two images per paper, i.e., two random figures, two random tables, or one random figure + table.
Models SN-F1 SN-Precision SN-Recall ITF-IDF
GPT-4o 47.73 42.09 55.48 5.95
w/ tables 46.76 41.32 54.17 5.53
w/ figures 46.62 41.20 54.04 5.48
w/ tables & figures 46.58 41.17 53.98 5.36
InternVL2-26B 41.91 41.02 43.28 1.48
w/ tables 40.55 40.37 42.91 1.46
w/ figures 42.88 42.10 43.76 1.46
w/ tables & figures 42.44 42.00 43.31 1.44
Our dataset covers both tables and figure illustrations extracted from the paper PDF as inputs. Intuitively, when reviewing a paper, both figures and tables are critical, not only for a better understanding, but also because some weaknesses are related to tables/figures. [^6] We find that there is approximately one human-written weakness related to figures or tables in each paper. Therefore, in Table 8, we adopt two LMMs to investigate the effectiveness of image inputs. Overall, image information, including both figures and tables, doesn’t bring significant performance improvement, i.e., only InternVL2 gains a performance boost after incorporating figures; while tables slightly drop both models’ results. This is probably because the LMMs cannot reason well over the information-intensive images, especially the table images.
5.4 ReviewCritique
Settings.
As individual review comments are split into multiple smaller segments (sentences), in order to avoid the performance variance that comes from the prompting, we follow Du et al. (2024) to utilize two prompting strategies. i) Labeling-All: given everything necessary including a list of indexed review segments, require the LLM to output a list of triples, like {id, reliable or not, explanation}. ii) Select-Deficient: Given everything necessary including a list of indexed review segments, require the LLM to output a list of tuples, {id, explanation}, when it believes the “id” corresponds to a deficient segment.
To further enhance evaluation robustness, we ensemble the results obtained from the two prompting strategies using two methods. i) Both “No”: if both prompts classify a segment as deficient, we consider it to be deficient. ii) Either “No”: if either of the prompts labels a segment as Deficient, we consider it to be deficient.
Main results.
Table 9: From (Du et al., 2024), various LLMs’ performances on the 11,376 instances of ReviewCritique. The best F1 score among different prompt methods for a single model is underlined. The best F1 score across all models is also bold.
Models Precision / Recall / F1
Labeling-All Select-Deficient Both “No” Either “No”
Open-source LLMs
Llama3-8B (AI@Meta, 2024) 7.73 / 45.95 / 12.22 11.47 / 30.29 / 14.88 11.37 / 21.27 / 12.46 8.19 / 53.61 / 13.35
Llama3-70B (AI@Meta, 2024) 13.63 / 42.49 / 18.19 13.95 / 31.16 / 17.46 16.16 / 23.51 / 16.67 12.46 / 50.02 / 18.43
Qwen2-72B (Bai et al., 2023) 9.97 / 26.60 / 12.96 11.35 / 34.61 / 14.64 9.07 / 15.13 / 9.62 10.49 / 43.00 / 15.16
Closed-source LLMs
Gemini 1.5 (Anil et al., 2023) 16.58 / 34.13 / 19.76 14.71 / 43.60 / 19.72 17.01 / 27.05 / 18.28 14.46 / 50.37 / 20.34
GPT-4 (OpenAI et al., 2023) 14.91 / 34.49 / 18.38 17.18 / 34.59 / 20.30 18.71 / 21.40 / 16.85 14.72 / 47.68 / 20.66
Claude Opus (Anthropic, 2024b) 16.86 / 34.26 / 20.35 17.69 / 26.61 / 18.71 17.14 / 18.70 / 15.78 16.94 / 42.12 / 21.99
We put the results of Du et al. (2024) in Table 9. Closed-source models (GPT-4, Claude Opus, and Gemini 1.5) generally outperform open-source models (Llama3-8B and 70B, Qwen2-72B) in F1 score. Claude Opus achieves the highest F1 scores, with GPT-4 and Gemini 1.5 performing slightly worse. Notably, recall scores are consistently higher than precision scores across all LLMs and prompting strategies, suggesting that LLMs tend to incorrectly identify segments as deficient. Despite the superior performance of the closed-source models, their F1 scores remain relatively low even with different prompt strategies, highlighting the challenges LLMs face in such expertise-intensive tasks and emphasizing the importance of human expertise in the meta-reviewing process.
$\mathcal{Q}$: How about the LLMs explanation quality regarding the deficient review?
As we also prompt the LLMs to generate the corresponding explanation on why they think each review segment is deficient, we report how well the model-generated deficiency explanation aligns with the human explanation.
Table 10: Evaluation of LLMs’ explanations for correctly identified deficient segments.
Model ROUGE-1/2/L/BERTScore
GPT-4 17.13 / 2.71 / 14.64 / 55.63
Claude Opus 20.18 / 3.69 / 17.52 / 57.28
Gemini 1.5 18.47 / 2.98 / 16.38 / 56.46
Llama3-8B 16.49 / 2.22 / 13.65 / 55.23
Llama3-70B 15.94 / 1.95 / 13.78 / 57.09
Qwen2-72B 17.07 / 3.00 / 14.69 / 56.88
We put the results of Du et al. (2024) in Table 10. The results in Table 10 show that overall scores for all LLMs are relatively low, indicating they can identify some Deficient segments but struggle to articulate their reasoning. Among the LLMs, Claude Opus achieves the highest scores across all metrics, suggesting its explanations align best with human annotators. Claude Opus also excels in identifying Deficient segments, as shown previously. GPT-4 and Gemini 1.5 show similar performance to Claude Opus. The open-source models, Llama3 (8B and 70B) and Qwen2-72B, generally score lower than the closed-source models.
6 Conclusion
In this work, we propose AAAR-1.0, a novel benchmark targeting a comprehensive evaluation of the current LLMs’ AI research capacity. AAAR-1.0 consists of distinct expertise-intensive tasks along with the curated evaluation metrics, and collect high-quality data by employing senior AI researchers. Extensive experiments highlight the challenges and values of AAAR-1.0.
Limitations
We shed light on two limitations of this work: i) As we gather data from open-source platforms such as arXiv and OpenReview, there is a possibility that current or future LLMs may be trained on the same source data utilized in our benchmark. This situation could influence the fairness of LLM comparisons and the conclusions drawn from this paper. While we acknowledge this potential data leakage, we maintain that our work can provide valuable insights and serve as an upper bound for some LLMs, particularly if they are indeed pretrained on those scientific papers. At present, this research can inform the design of current LLMs and agents, enhancing the community’s understanding of the strengths and limitations of using LLMs in scientific research. ii) Meanwhile, due to the high consumption of employing senior researchers in conducting data annotation, the data size for some tasks, such as ExpDesign, is relatively small. This might introduce variability in LLM performance, even with repeated runs. As this work marks the beginning of a series, we plan to release larger-scale datasets that will cover more diverse research tasks in AAAR-2.0.
Ethics Statement
Our study highlights that LLMs can be used to assist humans in AI research. This doesn’t mean we encourage the use of AI to replace human researchers. In contrast, we emphasize that the main responsibility for conducting scientific research should always lie with humans to prevent any societal risks, while the LLMs are only tools for making human research more efficient. To this end, we commit to careful distribution of the data collected in our research, ensuring it serves strictly for research purposes. Our goal is to mitigate the risks while maximizing the benefits offered by LLMs.
Acknowledgments
We would like to thank arXiv and OpenReview for releasing the paper source packages and the review comments. The authors would also like to thank Ibraheem Moosa and Sarkar Snigdha Sarathi Das for assisting in the data collection.
Appendices
Within this supplementary material, we elaborate on the following aspects:
  • Appendix A: Data Statistics and Diversity
  • Appendix B: Implementation Details
  • Appendix C: More Experiment Results and Details
  • Appendix D: Data Cases and Annotation Platform Illustration
  • Appendix E: Prompt Templates
Appendix A Data Statistics and Diversity
We provide the detailed data statistics of three datasets in our benchmark, as shown in Table 11, 12, and 13. We use the NLTK package [^7] https://www.nltk.org/ to tokenize words and count the length. When calculating the length of equations, we use the pylatexenc tool [^8] https://github.com/phfaist/pylatexenc to simplify the equations first.
Meanwhile, for the Weakness, we also plot the review scores distribution of the papers used in the dataset, as well as the track distribution. As can be found in Figure 5, our dataset has a decent distribution, where the papers are uniformly distributed across 13 tracks, and most papers’ scores ranged from 5 to 8 (i.e., most papers are weakly rejected or accepted).
[…]
# of classification instances 1,049
# of source papers 869
ave. “left” input context length (in words) 4,377
ave. “right” input context length (in words) 6,362
max “left” input context length (in words) 24,849
max “right” input context length (in words) 32,948
min “left” input context length (in words) 711
min “right” input context length (in words) 8
ave. “pos.” output equation length (in character) 55
ave. “neg.” output equation length (in character) 48
max “pos.” output equation length (in character) 1,039
max “neg.” output equation length (in character) 306
min “pos.” output equation length (in character) 6
min “neg.” output equation length (in character) 4
Table 12: The statistics of ExpDesign.
# of instances 100
# of source papers 100
ave. input context length (in words) 4,288
max input context length (in words) 9,799
min input context length (in words) 698
ave. # of input figures 2.6
max # of input figures 16.0
min # of input figures 0.0
ave. length of Experiment&Explanation list 5.7
ave. length per experiment (in words) 34.3
ave. length per explanation (in words) 27.1
max length of Experiment&Explanation list 13
max length per experiment (in words) 135
max length per explanation (in words) 89
min length of Experiment&Explanation list 2
min length per experiment (in words) 9
min length per explanation (in words) 9
Table 13: The statistics of Weakness.
# of instances 993
# of source papers 993
ave. input context length (in words) 9,811
max input context length (in words) 49,195
min input context length (in words) 24
ave. # of input figures 7.0
max # of input figures 37.0
min # of input figures 0.0
ave. # of input tables 4.3
max # of input tables 53.0
min # of input tables 0.0
ave. # of reviewers per paper 3.8
max # of reviewers per paper 9.0
min # of reviewers per paper 3.0
ave. # of weaknesses per reviewer 4.8
max # of weaknesses per reviewer 39.0
min # of weaknesses per reviewer 1.0
ave. length of weakness (in words) 39.1
max length of weakness (in words) 371.0
min length of weakness (in words) 2.0
(a) The review score distribution of the papers used in Weakness.
(b) The track distribution of the papers used in Weakness.
Figure 5: The data diversity illustration of Weakness, including the score distribution and track distribution of the papers used in our dataset.
Appendix B Implementation Details
B.1 Metric Details
When calculating the metrics, specifically for the similarity-based scores, we utilize SentenceBERT (Reimers, 2019) to encode each segment (e.g., each experiment idea in the list) into a dense vector, and then calculate the cosine similarity, [^9] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 which takes about 1GB of memory when running on a single A100 GPU.
B.2 LLMs Running Details
In our experiments, we utilize various LLMs, including both closed and open-sourced. We list the model weight sources for the open-source LLMs:
We use VLLM to unify the inference endpoints of all the above models. [^10] https://github.com/vllm-project/vllm We use Pytorch 2.4.0 with CUDA 12.1, and use 8 NVIDIA A100 GPUs for the LLMs inference.
Meanwhile, we use the gpt-4o-2024-08-06, gpt-4-1106-preview, o1-preview-2024-09-12, gemini-1.5-pro-002, and claude-3-5-sonnet-20240620 for the closed-source LLMs. We use LiteLLM to unify the API calling for all these LLMs. [^11] https://github.com/BerriAI/litellm
Given the unstable performance of LLMs, particularly closed-source ones, we run each model thrice during our experiments, selecting the median result from these repeated runs.
Appendix C More Experiment Results and Details
C.1 Human Evaluation on LLM-Generated Explanation
We ask 5 annotators to evaluate the LLM-generated explanations. Specifically, each of them is assigned 4 or 5 papers, along with the corresponding experiment lists. For each paper, the annotator is given 5 different models’ outputs (model names are anonymized), and the annotator has to decide if each LLM-generated explanation is acceptable according to the experiment. We show the human evaluation results in Table 4,
Appendix D Data cases and Annotation Platform Illustration
As shown in Figure 7, 8, and 9, we show the sample cases of the three tasks in AAAR-1.0. Meanwhile, we illustrate the screenshot of our annotation platform in Figure 6.
Figure 6: The annotation platform for collecting the annotation of ExpDesign. We ask annotators to first make comments on the Google Drive PDF, then move all the annotations to the online Google Doc (for further verification and discussion).
Appendix E Prompt Templates
In this appendix, we attach all the prompts used in this work, including prompts in data collection and model prediction, as shown in Figure 10, 11, and 12.
Figure 7: A sample case of EqInfer.
Figure 8: A sample case of ExpDesign.
Figure 9: A sample case of Weakness.
Figure 10: The prompts used in EqInfer, including both data collection and model prediction.
Figure 11: The prompts used in ExpDesign, including both data collection and model prediction.
Figure 12: The prompts used in Weakness.